# Time Series Anomaly Detection in the Frequency Domain with Statistical Reliability

### v1

@@ 3.4 Step 2: Refining Solution by Simulated Annealing
1:$\Delta E$ and $c$
2:Uniformly sample $\theta$ from $[0,1)$
3:if $\Delta E\leq 0$ then
@@
6:   if $\exp(-\Delta E/c)>\theta$ then
7:      status $\leftarrow$ Acceptance
8:   else
9:      status $\leftarrow$ Rejection
10:   end if
11:end if
12:status
Local search. In this paper, we consider four types of neighborhood operations applied to the current solution, i.e., adding, removing, and moving a CP (Lavielle, 1998) for a randomly selected frequency $d$, and merging two adjacent CP locations randomly selected from $\bm{\tau}$ at a random position between them. Schematic illustrations of these four operations are provided in Figures 3 and 4. We set the number of searching iterations at a specific temperature $c$ equal to the size of neighborhoods, i.e., $|\mathcal{D}^{\text{init}}|\cdot T$ (Aarts & Korst, 1989). Each operation is randomly selected from adding, removing, and moving one CP. Subsequently, two adjacent CP locations are merged only once because this operation significantly fluctuates the objective function value. When the operation is not possible (e.g., removing operation for a frequency with no CP), it is skipped. If no transition to neighborhoods occurs when these operations are repeated sufficiently at a certain temperature $c$, the search is terminated.
The setting of initial temperature. We determine the initial temperature $c_{0}$ by setting an acceptance ratio
$\chi(c_{0})=\frac{\#\text{ accepted transitions}}{\#\text{ proposed transitions}}$ (7)
to a desired value in a preliminary experiment of simulated annealing. In practice, we start by setting the temperature to a sufficiently small positive value, then multiply it with a constant factor $\lambda^{+}$, larger than 1, as follows
$c_{i+1}^{+}=\lambda^{+}c_{i}^{+},$
@@
Decrement of temperature. We employ a geometric cooling schedule, which is used as a practical method of temperature control, altough it does not guarantee the convergence to a global optimum. The decrement function of the $i$-th temperature $c_{i}$ is given as
@@
Algorithm 2 CP candidate selection using simulated annealing
1:$\bm{x}$
@@
4:$c\leftarrow c_{0}$ such that $\chi(c_{0})$ in (7) is the desired value
@@
7:      $d$ is randomly selected from $\mathcal{D}^{\text{init}}$
8:      The operation is randomly selected from adding, removing, and moving a CP for $d$
@@
10:      status $\leftarrow$ metropolis_algorithm($\Delta E(\mathcal{T}^{\prime},\mathcal{T},\bm{x}),c$)
@@
15:   Obtain $\bm{\mathcal{T}}^{\prime}$ by merging two adjacent CP locations in $\bm{\mathcal{T}}$
16:   status $\leftarrow$ metropolis_algorithm($\Delta E(\mathcal{T}^{\prime},\mathcal{T},\bm{x}),c$)
@@
23:   $c\leftarrow\lambda c$
@@
25:$\bm{\mathcal{T}}$
4 Selective Inference on CP candidate Locations
In this section, using the SI framework, we quantify the statistical significance of all locations selected as CP candidates by the algorithm $\mathcal{A}$ in Section 3 in the form of $p$-values. By setting the significance level $\alpha$ (e.g., 0.05 or 0.01) and considering CP candidate locations with the $p$-value $<\alpha$ as the final CPs, it is theoretically guaranteed that the false positive detection probabilities (type I error rates) of these final CPs is controlled below the specified significance level $\alpha$.
To formalize the SI framework, let us write the algorithm in Section 3 as
$\mathcal{A}:\bm{X}\mapsto\bm{\mathcal{T}},$
and the CP candidates obtained by applying the algorithm to the actual observed time series data $\bm{x}$ are expressed as
$\bm{\mathcal{T}}^{\rm det}=\left({\bm{\tau}^{\rm det}}^{(0)},\ldots,{\bm{\tau}^{\rm det}}^{(D-1)}\right)=\mathcal{A}(\bm{x}),$ (8)
where ${\bm{\tau}^{\rm det}}^{(d)}=\{{\tau_{1}^{\rm det}}^{(d)},\ldots,{\tau_{K^{(d)}}^{\rm det}}^{(d)}\}$ is the set of CP candidates for frequency $d\in\{0,\ldots,D-1\}$ 555 Note that, to ensure deterministic behavior of the algorithm $\mathcal{A}$, the random seed is fixed to a constant value at the beginning of the procedure.. Furthermore, let $\bm{\tau}^{\rm det}=\{\tau_{1}^{\rm det},\ldots,\tau_{K}^{\rm det}\}$ be the ordered set of CP candidate locations, $\mathcal{D}_{\tau_{k}^{\rm det}}\subseteq\{0,\dots,D-1\}$ be the set of frequencies that have CP candidates at a CP location $\tau_{k}^{\rm det}$, and $k^{(d)}\in[K^{(d)}]$ be the correponding index of the CP candidates for $d$ and $k\in[K]$, i.e., ${\tau_{k^{(d)}}^{\text{det}}}^{(d)}=\tau_{k}^{\text{det}}$ holds for all $d\in\mathcal{D}_{\tau_{k}^{\text{det}}}$ 666 These notations are somewhat intricate. For example, in the case of Figure 2, $D=5$ $\bm{\mathcal{T}}^{\rm det}=\left(\{97\},\{28\},\{28,70\},\{70\},\{\}\right)$ $\bm{\tau}^{\rm det}=\{28,70,97\}$ $K=3$ $\mathcal{D}_{\tau_{1}^{\rm det}}=\{d_{1},d_{2}\},\mathcal{D}_{\tau_{2}^{\rm det}}=\{d_{2},d_{3}\},\mathcal{D}_{\tau_{3}^{\rm det}}=\{d_{0}\}$ $k=3,1,1,2,2$ for $(d,k^{(d)})=(d_{0},1),(d_{1},1),(d_{2},1),(d_{2},2),(d_{3},1)$, respectively. .
4.1 Statistical Test on CP Locations
Hypotheses. For testing the statistical significance of CP location $\tau^{\rm det}_{k}$ for $k\in[K]$, we consider the following null hypothesis ${\rm H}_{0,k}$ and alternative hypothesis ${\rm H}_{1,k}$:
${\rm H}_{0,k}:\frac{1}{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}\,\sum_{t={\tau_{k^{(d)}-1}^{\rm det}}^{(d)}\,+1}^{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\mu_{t}^{(d)}$ $=\frac{1}{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\sum_{t={\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1}^{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}\mu_{t}^{(d)},~{}~{}~{}\forall d\in\mathcal{D}_{\tau_{k}^{\rm det}},$ (9)
@@ 4.1 Statistical Test on CP Locations
${\rm H}_{1,k}:\frac{1}{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}\,\sum_{t={\tau_{k^{(d)}-1}^{\rm det}}^{(d)}\,+1}^{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\mu_{t}^{(d)}$ $\neq\frac{1}{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\sum_{t={\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1}^{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}\mu_{t}^{(d)},~{}~{}~{}\exists d\in\mathcal{D}_{\tau_{k}^{\rm det}}.$ (10)
The above null hypothesis ${\rm H}_{0,k}$ indicates that, at the location $\tau_{k}^{\rm det}$, the means of $\mu_{t}^{(d)}$ in the left segment and the right segment are equal for all frequencies $d\in\mathcal{D}_{\tau_{k}^{\rm det}}$. On the other hand, the alternative hypothesis ${\rm H}_{1,k}$ implies that the means of the left and right segments are not equal in at least one of the frequencies.
Test statistic. For testing with the null hypothesis (9) and alternative hypothesis (10), we consider the test statistic constructed by computing the differences between averages of the complex spectra in the two segments before and after the CP location $\tau_{k}^{\text{det}}$ for each frequency $d\in\mathcal{D}_{\tau_{k}^{\text{det}}}$, and then aggregating their squared absolute values. Formally, the test statistic is defined as follows:
$T_{k}(\bm{X})=\sigma^{-1}\sqrt{\sum_{d\in\mathcal{D}_{\tau_{k}^{\text{det}}}}a_{k^{(d)}}\bigg{|}\bar{F}_{{{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}+1:{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}}^{(d)}-\bar{F}_{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1:{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}^{(d)}\,\bigg{|}^{2}},$ (11)
@@
$a_{k^{(d)}}=\frac{\left({{\tau_{k^{(d)}}^{\text{det}}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}\right)\left({\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}}^{\text{det}}}^{(d)}\right)c_{\text{sym}}^{(d)}}{\left({{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}\right)M}\in\mathbb{R}$
is required for scaling the test statistic, and the spectral averages in the two segments before and after the CP location $\tau_{k}^{\text{det}}$ for frequency $d\in\mathcal{D}_{\tau_{k}^{\text{det}}}$ are respectively denoted as
$\bar{F}_{{{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}+1:{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}}^{(d)}$ $={\frac{1}{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}}\,\sum_{t={\tau_{k^{(d)}-1}^{\rm det}}^{(d)}+1}^{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}F_{t}^{(d)}$
$={\frac{1}{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}}\left(\bm{1}_{{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}+1:{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\otimes\bm{w}_{M}^{(d)}\right)^{\top}\bm{X}\in\mathbb{C},$
$\bar{F}_{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1:{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}^{(d)}$ $=\frac{1}{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\sum_{t={\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1}^{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}F_{t}^{(d)}$
$=\frac{1}{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\left(\bm{1}_{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1:{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}\otimes\bm{w}_{M}^{(d)}\right)^{\top}\bm{X}\in\mathbb{C}.$
By introducing a projection matrix defined as
$P_{k}=\sum_{d\in\mathcal{D}_{\tau_{k}^{\text{det}}}}a_{k^{(d)}}{\bm{v}}_{k^{(d)}}{{\bm{v}}^{\ast}}_{k^{(d)}}^{\top}\in\mathbb{R}^{N\times N},$
the test statistic in (11) is simply written as
$T_{k}(\bm{X})=\sigma^{-1}||P_{k}\bm{X}||,$ (12)
where
${\bm{v}}_{k^{(d)}}={\frac{1}{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}-{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}}}\left(\bm{1}_{{\tau_{k^{(d)}-1}^{\rm det}}^{(d)}+1:{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\otimes\bm{w}_{M}^{(d)}\right)-\frac{1}{{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}-{\tau_{k^{(d)}}^{\text{det}}}^{(d)}}\left(\bm{1}_{{\tau_{k^{(d)}}^{\text{det}}}^{(d)}+1:{\tau_{k^{(d)}+1}^{\rm det}}^{(d)}}\otimes\bm{w}_{M}^{(d)}\right)\in\mathbb{C}^{N},$
and ${{\bm{v}}^{\ast}}_{k^{(d)}}$ is the complex conjugate of ${\bm{v}}_{k^{(d)}}$.
Naive $p$-value. The sampling distribution of the test statistic in (11) and (12) is highly complicated. However, if we “forget” the fact that the CP candidates are selected by looking at the sequence $\bm{X}$, the matrix $P_{k}$ in (12) does not depend on $\bm{X}$, meaning that the test statistic $T_{k}(\bm{X})$ simply follows a $\chi$-distribution with ${\rm tr}(P_{k})$ degrees of freedom under the null hypothesis. The $p$-values obtained by this sampling distribution are referred to as naive $p$-values and computed as
$p_{k}^{\text{naive}}=\mathbb{P}_{\text{H}_{0},k}(T_{k}(\bm{X})\geq T_{k}(\bm{x})).$
Of course, since the CP candidates are selected based on the sequence $\bm{X}$, it is not appropriate to quantify the statistical significance using naive $p$-values. If naive $p$-values are mistakenly used, the type I error rate can become significantly larger than the significance level $\alpha$.
4.2 Computing Selective $p$-values
Selective Inference (SI). To address the above issue, we consider the sampling distribution of the test statistic conditional on the event that the detected CP candidates $\bm{\mathcal{T}}$ for a random sequence $\bm{X}$ is the same as $\bm{\mathcal{T}}^{\text{det}}$ for the observed sequence $\bm{x}$, that is,
$T_{k}(\bm{X})\,|\,\{\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x})\}.$ (13)

### v2

@@ 3.4 Step 2: Refining the Solution by Simulated Annealing
1:Difference of objective function values $\Delta E$, temperature $c$
2:Uniformly sample $\xi$ from $[0,1)$
3:if $\Delta E+c\ln\xi<0$ then
@@
6:   status $\leftarrow$ Rejection
7:end if
8:status
Local search.
In this paper, we consider four types of neighborhood operations applied to the current solution, i.e., adding, removing, and moving a CP (Lavielle, 1998) for a uniformly selected frequency $d$, and merging two adjacent CP locations uniformly sampled from $\bm{\tau}$ at a random position between them. Schematic illustrations of these four operations are provided in Figures 2 and 3. We set the number of searching iterations at a specific temperature $c$ equal to the size of neighborhoods, i.e., $|\mathcal{D}^{\text{init}}|\cdot T$ (Aarts & Korst, 1989). Each operation is uniformly selected from adding, removing, and moving one CP. Subsequently, two adjacent CP locations are merged only once because this operation significantly fluctuates the objective function value. When the operation is not possible (e.g., removing operation for a frequency with no CP), it is skipped. If no transition to neighborhoods occurs when these operations are repeated sufficiently at a certain temperature $c$, the search is terminated.
@@
Setting of initial temperature.
We determine an initial temperature $c_{0}$ by setting an acceptance ratio
$\eta(c_{0})=\frac{\#\text{ accepted transitions}}{\#\text{ proposed transitions}}$
to a desired value in a preliminary experiment of simulated annealing. In practice, we start by setting the temperature to a sufficiently small positive value, then multiply it with a constant factor $\lambda^{+}$, larger than 1, as follows:
$c_{i+1}^{+}=\lambda^{+}c_{i}^{+},$
Decrement of temperature.
We employ a geometric cooling schedule, which is used as a practical method of temperature control, altough it does not guarantee the convergence to a global optimum. The decrement function of the $i$-th temperature $c_{i}$ is given as
@@
Algorithm 2 CP Candidate Selection by Simulated Annealing
1:Time sequence $\bm{x}$
@@
4:Determine the initial temperature $c$ in a preliminary experiment
@@
7:      Frequency $d$ is uniformly sampled from $\mathcal{D}^{\text{init}}$
8:      The operation is uniformly sampled from adding, removing, and moving a CP for $d$
@@
10:      status $\leftarrow$ metropolis_algorithm($\Delta E(\bm{\mathcal{T}}^{\prime},\bm{\mathcal{T}},\bm{x}),c$)
@@
15:   Obtain $\bm{\mathcal{T}}^{\prime}$ by merging uniformly sampled two adjacent CP locations in $\bm{\mathcal{T}}$
16:   status $\leftarrow$ metropolis_algorithm($\Delta E(\bm{\mathcal{T}}^{\prime},\bm{\mathcal{T}},\bm{x}),c$)
@@
23:   Update the temperature using a constant factor $\lambda$ as $c\leftarrow\lambda c$
@@
25:CP candidates $\bm{\mathcal{T}}$
4 Selective Inference on CP Candidate Location
In this section, using the SI framework, we quantify the statistical significance of all locations selected as CP candidates by the algorithm in Section 3 in the form of $p$-values. By setting the significance level $\alpha$ (e.g., 0.05 or 0.01) and considering CP candidate locations with the $p$-values below $\alpha$ as the final CPs, it is theoretically guaranteed that the false positive detection probabilities (type I error rates) of the final CPs are controlled below the specified significance level $\alpha$.
To formalize the SI framework, let us write the CP candidate selection algorithm in Section 3 as $\mathcal{A}\colon\bm{X}\mapsto\bm{\mathcal{T}}$ 555 Note that, to ensure deterministic behavior of the algorithm $\mathcal{A}$, the random seed is fixed to a constant value at the beginning of the procedure.. As detailed in Section 3, the output of this algorithm is a collection of CP candidates represented as $\bm{\mathcal{T}}=(\bm{\tau}^{(0)},\ldots,\bm{\tau}^{(D-1)})$, where each $\bm{\tau}^{(d)}=\{\tau_{1}^{(d)},\ldots,\tau_{K^{(d)}}^{(d)}\}\subseteq[T-1]$ is the ordered set of CP candidates for frequency $d\in\{0,\ldots,D-1\}$. The final set of selected CP candidate locations $\bm{\tau}$ is obtained by taking the union of these sets of CP candidates, i.e., $\bm{\tau}=\bigcup_{d=0}^{D-1}\bm{\tau}^{(d)}=\{\tau_{1},\ldots,\tau_{K}\}\subseteq[T-1]$. Subsequently, we quantify the statistical significance of each CP candidate location $\tau_{k}\in\bm{\tau}$ for $k\in[K]$, in the form of $p$-values within the SI framework. Therefore, without loss of generality, we formulate the hypothesis testing problem for the $k$-th CP candidate location $\tau_{k}$ for simplicity.
4.1 Statistical Test on CP Candidate Location
Hypotheses.
To test the statistical significance of CP candidate location $\tau_{k}$, we consider whether a mean-shift occurs before and after this location $\tau_{k}$ at any frequency for which this location was selected as a CP candidate. To formalize this, let $\mathcal{D}_{k}$ denote the set of frequencies where $\tau_{k}$ is selected as a CP candidate location, i.e., $\mathcal{D}_{k}=\{d\in\{0,\ldots,D-1\}\mid\tau_{k}\in\bm{\tau}^{(d)}\}$. For each $d\in\mathcal{D}_{k}$, let $\tau_{\mathrm{pre}}^{(d)}$ and $\tau_{\mathrm{suc}}^{(d)}$ represent the CP candidate locations immediately before and after $\tau_{k}$ at frequency $d$, respectively. Note that although $\tau_{\mathrm{pre}}^{(d)}$ and $\tau_{\mathrm{suc}}^{(d)}$ depend on $\tau_{k}$, this dependence is omitted from the notation for simplicity. Given these notations, we formulate the null hypothesis $\mathrm{H}_{0,k}$ and alternative hypothesis $\mathrm{H}_{1,k}$ as follows:
$\mathrm{H}_{0,k}\colon\frac{1}{\tau_{k}-\tau_{\mathrm{pre}}^{(d)}}\sum_{t=\tau_{\mathrm{pre}}^{(d)}+1}^{\tau_{k}}\mu_{t}^{(d)}=\frac{1}{\tau_{\mathrm{suc}}^{(d)}-\tau_{k}}\sum_{t=\tau_{k}+1}^{\tau_{\mathrm{suc}}^{(d)}}\mu_{t}^{(d)},\ \forall d\in\mathcal{D}_{k},$ (8)
@@ 4.1 Statistical Test on CP Candidate Location
$\mathrm{H}_{1,k}\colon\frac{1}{\tau_{k}-\tau_{\mathrm{pre}}^{(d)}}\sum_{t=\tau_{\mathrm{pre}}^{(d)}+1}^{\tau_{k}}\mu_{t}^{(d)}\neq\frac{1}{\tau_{\mathrm{suc}}^{(d)}-\tau_{k}}\sum_{t=\tau_{k}+1}^{\tau_{\mathrm{suc}}^{(d)}}\mu_{t}^{(d)},\ \exists d\in\mathcal{D}_{k}.$ (9)
Test statistic.
We define the test statistic to test with the null hypothesis (8) and alternative hypothesis (9) as follows. First, for each frequency $d\in\mathcal{D}_{k}$, we consider the difference between the average values of the complex spectra in the two segments before and after the location $\tau_{k}$. Then, we aggregate these differences across all frequencies $d\in\mathcal{D}_{k}$ to define the test statistic. Therefore, the test statistic $T_{k}(\bm{X})$ is formulated as
$T_{k}(\bm{X})=\sigma^{-1}\sqrt{\sum_{d\in\mathcal{D}_{k}}\frac{a_{\mathrm{len}}^{(d)}c_{\mathrm{sym}}^{(d)}}{M}\left|\bar{F}_{\tau_{\mathrm{pre}}^{(d)}+1:\tau_{k}}^{(d)}-\bar{F}_{\tau_{k}+1:\tau_{\mathrm{suc}}^{(d)}}^{(d)}\right|^{2}},$ (10)
where $a_{\mathrm{len}}^{(d)}=(\tau_{\mathrm{suc}}^{(d)}-\tau_{k})(\tau_{k}-\tau_{\mathrm{pre}}^{(d)})/(\tau_{\mathrm{suc}}^{(d)}-\tau_{\mathrm{pre}}^{(d)})\in\mathbb{R}$ is a scaling factor for correcting the segment lengths before and after the location $\tau_{k}$ for each frequency $d\in\mathcal{D}_{k}$, $c_{\mathrm{sym}}^{(d)}\in\mathbb{R}$ defined in (6) is required for considering the complex conjugate symmetry of the spectrum, and $\bar{F}_{s:e}^{(d)}\in\mathbb{C}$ is the average value of the complex spectra in the segment from time points $s$ to $e$ for frequency $d\in\mathcal{D}_{k}$, i.e.,
$\bar{F}_{s:e}^{(d)}=\frac{1}{e-s+1}\sum_{t=s}^{e}F_{t}^{(d)}=\frac{1}{e-s+1}\left(\bm{1}_{s:e}\otimes\bm{w}_{M}^{(d)}\right)^{\top}\bm{X}.$ (11)
An important point when conducting the statistical test within the SI framework is that the test statistic is represented as the norm of a projection of the data, i.e., $T_{k}(\bm{X})=\sigma^{-1}||P_{k}\bm{X}||$, where $P_{k}\in\mathbb{R}^{N\times N}$ is a projection matrix that depends on the observed time series $\bm{x}$ only through the selected CP candidates $\mathcal{A}(\bm{x})$ (see the next subsection for further details). We now define the orthogonal projection matrix $P_{k}$ as
$P_{k}=\sum_{d\in\mathcal{D}_{k}}\frac{a_{\mathrm{len}}^{(d)}c_{\mathrm{sym}}^{(d)}}{M}\bm{v}^{(d)}\bm{v}^{(d)\ast\top}\in\mathbb{R}^{N\times N},$ (12)
@@
$\bm{v}^{(d)}=\frac{1}{\tau_{k}-\tau_{\mathrm{pre}}^{(d)}}\left(\bm{1}_{\tau_{\mathrm{pre}}^{(d)}+1:\tau_{k}}\otimes\bm{w}_{M}^{(d)}\right)-\frac{1}{\tau_{\mathrm{suc}}^{(d)}-\tau_{k}}\left(\bm{1}_{\tau_{k}+1:\tau_{\mathrm{suc}}^{(d)}}\otimes\bm{w}_{M}^{(d)}\right)\in\mathbb{C}^{N}$ (13)
is a vector that depends on the selected CP candidates $\mathcal{A}(\bm{x})$, and $\bm{v}^{(d)\ast}$ is the complex conjugate of $\bm{v}^{(d)}$. Using (10)–(13), the test statistic $T_{k}(\bm{X})$ can be rewritten as
$T_{k}(\bm{X})$ $=\sigma^{-1}\sqrt{\sum_{d\in\mathcal{D}_{k}}\frac{a_{\mathrm{len}}^{(d)}c_{\mathrm{sym}}^{(d)}}{M}|\bm{v}^{(d)\top}\bm{X}|^{2}}$
$=\sigma^{-1}\sqrt{\bm{X}^{\top}\left(\sum_{d\in\mathcal{D}_{k}}\frac{a_{\mathrm{len}}^{(d)}c_{\mathrm{sym}}^{(d)}}{M}\bm{v}^{(d)}\bm{v}^{(d)\ast\top}\right)\bm{X}}$
$=\sigma^{-1}\sqrt{\bm{X}^{\top}P_{k}\bm{X}}=\sigma^{-1}||P_{k}\bm{X}||~{}~{}~{}(\because P_{k}=P_{k}^{2},P_{k}=P_{k}^{\top}).$ (14)
4.2 Computing Selective $p$-value
Selective Inference (SI).
To compute the $p$-value, we need to identify the sampling distribution of the test statistic $T_{k}(\bm{X})$. However, as the projection matrix $P_{k}$ within the test statistic $T_{k}(\bm{X})$ depends on the CP candidates $\mathcal{A}(\bm{X})$ (which, in turn, depends on the sequence $\bm{X}$ through the CP candidate selection algorithm $\mathcal{A}$), the sampling distribution of the test statistic $T_{k}(\bm{X})$ is too complicated to characterize. Within the SI framework, to address this challenge, we consider the sampling distribution of the test statistic conditional on the event that the selected CP candidates $\mathcal{A}(\bm{X})$ for a random sequence $\bm{X}$ is the same as $\mathcal{A}(\bm{x})$ for the observed sequence $\bm{x}$, that is,
$T_{k}(\bm{X})\mid\{\mathcal{A}(\bm{X})=\mathcal{A}(\bm{x})\}.$ (15)